# SAGE: Spatial-visual Adaptive Graph Exploration for Visual Place Recognition

### v2

@@ 4.2 Implementation Details
SAGE is built upon the EMVP framework (emvp), which we reproduced from its publication for a fair comparison as the official code is unavailable. We fine-tune two Vision Transformer backbones, ViT-B and ViT-L, which we denote as SAGE-B and SAGE-L, respectively. The Feature Compression ($\mathcal{F}_{C}$) and Feature Probing ($\mathcal{F}_{P}$) branches are implemented as two-layer MLPs, reducing the feature dimensions to $D=128$ and $K=64$, respectively. The InteractHead module is implemented as a two-layer Transformer encoder with a model dimension of 768, 16 attention heads, and a feed forward network with dimension 1024. During training we freeze the backbone and adopt DPN for PEFT, which adaptively preserves task-specific information while greatly reducing the number of trainable parameters. Each training mini-batch is constructed with equal contributions from MSLS (all non-panoramic images from the training set) and GSV-Cities (0.56M images from 67K places). For every batch we build a new sparse geo-visual graph, and we sample $P=15$ sequences from the same city and recompute cliques dynamically during training. The thresholds for this process are set to $\tau_{1}=25$ m and $\tau_{2}=-2.88\times 10^{3}$. Input images are resized to $224\times 224$ during training and $322\times 322$ during inference. All experiments are implemented in PyTorch and run efficiently on a single NVIDIA A100 GPU. We fine-tune models for 10 epochs and select the checkpoint with the highest Recall@1 on Pitts30k-test for evaluation on the other benchmarks.
@@ 4.3 Comparisons with SOTA Methods
In this section we provide a comprehensive comparison between our proposed SAGE and a range of state-of-the-art VPR methods. The comparison includes: NetVLAD (netvlad), SFRS (sfrs), CosPlace (cosplace), MixVPR (mixvpr), EigenPlaces (eigenplaces), CricaVPR (cricavpr), SALAD (salad), BOQ (boq), SALAD-CM (salad-cm), SuperVLAD (supervlad), EMVP (emvp), and the two re-ranking (two-stage) pipelines SelaVPR (selavpr) and FoL (FoL). Re-ranking pipelines add a computationally intensive local feature matching stage. EMVP and FoL are the current SOTA for single-stage global retrieval and two-stage VPR pipelines, respectively; implementation details for compared methods are given in App. A.4.
@@ 4.4 Ablation Study
Adding GWS to a SoftP configuration produces modest and unstable gains, suggesting sampling alone cannot exploit graph dynamics without Online Graph Creation. Enabling OGC and GWS while retaining CFP yields a notable improvement, for example Pitts30k-test R@1 of 94.9%, but still falls short of the configuration that includes SoftP. The full SAGE configuration achieves the best results, with R@1 of 95.4% on Pitts30k-test and R@1 of 98.0% on SPED, surpassing the baseline and intermediate variants. The greedy weighted clique expansion complements SoftP and OGC by focusing training on the most informative clusters and enhancing descriptor discriminability. As detailed in Tab. 6, we conducted an ablation study on the internal dimensions of the InteractHead to determine its optimal configuration.
($d_{\text{model}},d_{\text{ff}}$) SPED Pitts30k
R@1 R@5 R@10 R@1 R@5 R@10
(512, 1024) 98.8 99.5 99.7 95.5 97.6 98.2
(768, 1536) 98.6 99.3 99.7 96.0 97.9 98.4
(768, 1024) 98.9 99.7 100 95.8 97.8 98.4
To evaluate the computational cost of our dynamic sampling strategy, we conducted an ablation study comparing the runtime and performance of online versus offline graph creation. As shown in Table 7, our online approach incurs a modest 17.7% increase in per-epoch training time. However, this modest overhead is a worthwhile investment, as it translates directly to superior accuracy, with the online SAGE model achieving higher recall on both SPED and MSLS-val.
@@
To demonstrate SAGE’s learning efficiency, we tracked its early-stage training performance on MSLS-val against a baseline using the Cliquemining (CM) strategy. As shown in Table 8, our model establishes a clear advantage by the fourth epoch (93.4% vs. 92.7% at R@1). The widening performance gap confirms that our dynamic sampling strategy fosters more efficient learning, leading to superior performance within the same epochs.
@@ 5 Conclusion
We introduce SAGE, a novel framework for Visual Place Recognition that excels under drastic appearance and viewpoint variations. SAGE strategically enhances feature discriminability with a Soft Probing module and an InteractHead, while simultaneously optimizing the training process through online geo-visual graph construction and greedy weighted sampling for dynamic hard-negative mining. Extensive experiments on eight diverse benchmarks demonstrate that SAGE sets a new state-of-the-art while maintaining high parameter efficiency. We acknowledge a performance drop in highly dynamic environments. Future work will primarily focus on tackling even more extreme domain shifts and exploring the fusion of multi-modal cues like semantics, to further improve robustness.
@@ A.1 Visualization of results
Figure 9: Visualization of SoftP’s learned feature importance. Each pair consists of an input image and its corresponding heatmap, where warmer colors indicate regions assigned higher importance by SoftP. The visualizations demonstrate that SoftP automatically learns to focus on stable, fine-grained landmarks (e.g., building facades, structural details) while effectively ignoring non-discriminative regions (sky, road) and transient objects (vehicles, pedestrians).
@@ A.2 Additional Results
Building on this, Tab. 10 extends the evaluation to more demanding datasets characterized by severe domain shifts, including AmsterTime (historical vs. modern), Tokyo24/7 (extreme viewpoint and time-of-day changes), Pitts250k-test (large-scale urban scenes), and Eynsham (rural route). The performance trend remains robust, with accuracy scaling gracefully with descriptor dimensionality. The improvement is particularly pronounced on AmsterTime, where the R@1 score surges from 23.4% at 128-D to 83.5% at 8448-D. This demonstrates SAGE’s exceptional capability to handle extreme appearance variations, validating the effectiveness of our proposed dynamic geo-visual graph exploration and feature enhancement strategies.
@@
To further assess the generalization capabilities of our method, we conducted experiments on the SF-small benchmark. Derived from SF-XL (cosplace), this dataset is particularly well-suited for evaluating generalization as its queries are drawn from distinct, non-adjacent locations.
@@
To further validate the effectiveness of SAGE, we present an extended comparison against several recently published SOTA methods from top-tier venues. These include VLAD-BuFF (VLAD-BuFF), EDTFormer (jin2025edtformer), and EffoVPR (tzachor2024effovpr). The results, shown in Table 13, demonstrate that SAGE consistently outperforms these strong baselines. SAGE achieves superior performance across the majority of benchmarks. Although EffoVPR obtains a slightly higher R@1 on Tokyo24/7, it is a two-stage reranking method, which introduces additional computational overhead.
@@ A.3 Dataset Details
Pitts30k-test (pitts). The Pitts30k-test dataset is a subset of the Pittsburgh 250k dataset collected from Google Street View panoramas with GPS labels. It contains urban street-view images from Pittsburgh, Pennsylvania, USA, covering diverse city environments such as roads, bridges, and buildings, with large variations in viewpoint, season, and illumination. Each location provides multiple viewpoint images, and the dataset is primarily used for evaluating VPR models in urban scenarios.
@@
SPED (SPED). The SPED dataset, also known as the Specific PlacEs Dataset, consists of images captured by fixed surveillance cameras over extended time periods, covering significant changes in illumination, weather, and seasons. It contains unique locations, each with hundreds of images taken at different times of day and year, providing a challenging benchmark for long-term VPR under extreme appearance variations.
@@
MSLS-val (msls). The MSLS-val dataset is the validation split of the Mapillary Street-Level Sequences dataset, which contains street-level imagery sequences from cities worldwide captured with various devices such as smartphones, dashcams, and professional mapping cameras. The validation set covers multiple cities, seasons, and weather conditions, and is widely used for tuning and evaluating VPR models under diverse geographic and environmental variations.
@@
Nordland (nordland). The Nordland dataset consists of front-facing video recordings from a train journey along the same railway route in Norway, captured in all four seasons: spring, summer, autumn, and winter. The route and camera viewpoints are fixed, ensuring identical spatial structure while presenting extreme appearance changes due solely to seasonal and weather differences, making it a valuable benchmark for cross-season VPR research.
@@
Tokyo247 (tokyo247). The Tokyo247 dataset contains urban street-view images from Tokyo, Japan, primarily sourced from Google Street View panoramas with GPS labels, and is named for including both day and night imagery to represent long-term scene changes. It features dense metropolitan areas with tall buildings as well as some open spaces, offering large variations in viewpoint, illumination, and dynamic objects, and is used to evaluate VPR models in complex urban environments.
@@
AmsterTime (amstertime). The AmsterTime dataset is a time-lapse street-view dataset of Amsterdam, containing images of the same locations captured in different years, such as 2008 and 2014, using Google Street View panoramas. It reflects long-term changes in building facades and urban infrastructure, providing a benchmark for evaluating VPR models under temporal urban transformations.
@@
Pitts250k-test (pitts). The Pitts250k-test dataset is the test split of the Pittsburgh 250k dataset, consisting of street-view images from Pittsburgh and surrounding areas, captured from Google Street View panoramas with GPS labels. It offers diverse urban scenes with dense road networks, buildings, and bridges, and serves as a large-scale benchmark for evaluating the scalability and robustness of VPR systems.
@@
Eynsham (eynsham; berton2022deep). The Eynsham dataset is a GPS-synchronized street-view image sequence collected along approximately 35 km of driving routes in Eynsham, Oxfordshire, UK, using a vehicle-mounted camera. It provides continuous video frames with precise ground-truth positions, making it suitable for sequence-based place recognition and loop closure detection experiments.
@@
SF-XL (cosplace). The San Francisco eXtra Large (SF-XL) dataset is a massive and dense new benchmark designed to push visual geo-localization research towards realistic, city-wide applications. Comprising over 41 million images captured across a decade, it presents significant real-world challenges such as long-term temporal variations and a domain shift between its Street View database and crowd-sourced queries.
@@ A.4 Compared Methods Details
NetVLAD (netvlad)https://github.com/Nanne/pytorch-NetVlad. A classic VPR method with a learnable VLAD layer pluggable into any CNN. Uses VGG-16 backbone, trained on Pitts30k, optimized via weakly supervised ranking loss, outperforming traditional methods.
@@
SFRS (sfrs)https://github.com/yxgeee/OpenIBL. Addresses GPS noise by mining hard positives via self-supervised fine-grained region similarities, multi-generation training. Based on NetVLAD, VGG-16 backbone, trained on Pitts30k, outperforming state-of-the-art then.
@@
CosPlace (cosplace)https://github.com/gmberton/CosPlace. Solves scalability in large-scale localization by framing training as classification. Constructs SF-XL dataset, uses CosPlace Groups. VGG-16/ResNet backbone, outputs 512D descriptors, low memory usage, suitable for city-scale applications.
@@
MixVPR (mixvpr)https://github.com/amaralibey/MixVPR. This is a novel holistic feature aggregation method for VPR. It takes feature maps from pre-trained backbones as global features and iteratively incorporates global relationships into each feature map through stacked Feature-Mixer blocks (composed solely of multi-layer perceptrons), without the need for local or pyramidal aggregation. Using backbones like ResNet and trained on datasets such as GSV-Cities, it outperforms existing methods by a large margin with less than half the number of parameters compared to CosPlace and NetVLAD.
@@
R2Former (r2former)https://github.com/bytedance/R2Former. This method introduces R2Former, a unified framework that employs a pure Transformer architecture to handle both global retrieval and local reranking in a single, end-to-end model. Its novel reranking module replaces slow geometric verification with a learnable Transformer that analyzes richer cues like feature correlation and attention, achieving state-of-the-art accuracy while dramatically reducing inference time and memory usage.
@@
EigenPlaces (eigenplaces)https: //github.com/gmberton/EigenPlaces. This method embeds viewpoint robustness into learned global descriptors by clustering training data to explicitly present the model with different views of the same points of interest, without extra supervision. Using backbones like VGG-16 or ResNet with GeM pooling and trained on the SF-XL dataset, it outperforms state-of-the-art methods on most datasets, requiring $60\%$ less GPU memory for training and using $50\%$ smaller descriptors.
@@
SelaVPR (selavpr)https://github.com/Lu-Feng/SelaVPR. It proposes a hybrid global-local adaptation method that adapts pre-trained foundation models (e.g., DINOv2) via lightweight adapters without modifying the pre-trained model parameters, efficiently generating global features for candidate retrieval and local features for re-ranking. It also introduces a mutual nearest neighbor local feature loss to avoid time-consuming spatial verification. Outperforming state-of-the-art methods on benchmarks like MSLS, it consumes only about $3\%$ of the retrieval time of RANSAC-based two-stage methods.
@@
CricaVPR (cricavpr)https://github.com/Lu-Feng/CricaVPR. It proposes cross-image correlation-aware representation learning, using attention to correlate features of multiple images in a batch, enabling each image feature to gain useful information from others for enhanced robustness. A multi-scale convolution-enhanced adaptation method is designed to insert lightweight adapters into frozen pre-trained foundation models (e.g., DINOv2) to introduce multi-scale local information. It outperforms state-of-the-art methods by a large margin on multiple benchmarks with shorter training time.
@@
SALAD (salad)https://github.com/serizba/salad. It reformulates NetVLAD’s soft assignment of local features to clusters as an optimal transport problem, considering both feature-to-cluster and cluster-to-feature relations, and introduces a ’dustbin’ cluster to discard non-informative features. Using DINOv2 as the backbone with fine-tuning, it trains in only 4 epochs. This single-stage method outperforms both single-stage and two-stage methods, with fast inference speed.
@@
BoQ (boq)https://github.com/amaralibey/ Bag-of-Queries. This method learns a set of global queries and uses cross-attention to probe input features for consistent information aggregation. It supports both CNN and Vision Transformer backbones, trained on the GSV-Cities dataset. As a one-stage global retrieval method without re-ranking, surpasses two-stage methods, and is fast and efficient.
@@
SALAD-CM (salad-cm)https://github.com/serizba/cliquemining.This work addresses the insufficient Geographic Distance Sensitivity (GDS) of existing VPR models by proposing a novel sample mining strategy. It constructs a graph of visually similar images and samples cliques (sets of geographically close images) from the graph as training batches to enhance the model’s ability to distinguish small-range geographic distances. Based on models like DINOv2 SALAD and MixVPR, trained on densely sampled datasets such as MSLS and Nordland, it significantly improves recall without increasing inference computational overhead.
@@
SuperVLAD (supervlad)https://github.com/lu-feng/SuperVLAD. This method improves NetVLAD by removing cluster centers and using a small number of clusters , enhancing cross-domain generalization and simplifying the model. It also proposes 1-Cluster VLAD, which generates extremely low-dimensional descriptors by introducing “ghost clusters” and outperforms methods like GeM pooling with the same dimension. Using Transformer backbones (e.g., DINOv2) and trained on datasets like Pitts30k, it outperforms existing methods with lower feature dimensions.
@@
EMVP (emvp)https://github.com/vincentqqb/EMVP. This method leverages Visual Foundation Models(e.g., DINOv2) and proposes a Centroid-Free Probing (CFP) stage that uses second-order features to better adapt VFM descriptors. It introduces a Dynamic Power Normalization (DPN) module to adaptively preserve task-specific information in both recalibration and CFP stages, forming a Parameter Efficiency Fine-Tuning (PEFT) pipeline. It achieves excellent performance on datasets like MSLS and Pitts250k, saving $64.3\%$ trainable parameters compared to existing state-of-the-art PEFT methods.
@@
FoL (FoL)https://github.com/chenshunpeng/FoL. This two-stage VPR method models reliable discriminative regions via Extraction-Aggregation Spatial Alignment Loss (SAL) and Foreground-Background Contrast Enhancement Loss (CEL), guiding global feature generation and efficient re-ranking. It introduces a weakly supervised local feature training strategy based on pseudo-correspondences and a discriminative region-guided efficient re-ranking pipeline (xu2024local). Using DINOv2 as the backbone and trained on GSV-Cities, it outperforms existing two-stage methods on multiple benchmarks with higher computational efficiency.
@@
VLAD-BuFF (VLAD-BuFF)https://github.com/Ahmedest61/VLAD-BuFF. This method improves visual place recognition accuracy by implementing a burst-aware weighting mechanism that discounts repetitive features to emphasize more distinctive visual cues. Simultaneously, it achieves high computational efficiency by using a pre-projection layer for local features, initialized with PCA, which enables rapid aggregation in a lower-dimensional space while maintaining high recall.

### v3

@@ 4.2 Implementation Details
SAGE is built upon the EMVP framework (Qiu et al., 2024), which we reproduced from its publication for a fair comparison as the official code is unavailable. We fine-tune two Vision Transformer backbones, ViT-B and ViT-L, which we denote as SAGE-B and SAGE-L, respectively. The Feature Compression ($\mathcal{F}_{C}$) and Feature Probing ($\mathcal{F}_{P}$) branches are implemented as two-layer MLPs, reducing the feature dimensions to $D=128$ and $K=64$, respectively. The InteractHead module is implemented as a two-layer Transformer encoder with a model dimension of 768, 16 attention heads, and a feed forward network with dimension 1024. During training we freeze the backbone and adopt DPN for PEFT, which adaptively preserves task-specific information while greatly reducing the number of trainable parameters. Each training mini-batch is constructed with equal contributions from MSLS (all non-panoramic images from the training set) and GSV-Cities (0.56M images from 67K places). For every batch we build a new sparse geo-visual graph, and we sample $P=15$ sequences from the same city and recompute cliques dynamically during training. The thresholds for this process are set to $\tau_{1}=25$ m and $\tau_{2}=-2.88\times 10^{3}$. Input images are resized to $224\times 224$ during training and $322\times 322$ during inference. All experiments are implemented in PyTorch and run efficiently on a single NVIDIA A100 GPU. We fine-tune models for 10 epochs and select the checkpoint with the highest Recall@1 on Pitts30k-test for evaluation on the other benchmarks.
@@ 4.3 Comparisons with SOTA Methods
In this section we provide a comprehensive comparison between our proposed SAGE and a range of state-of-the-art VPR methods. The comparison includes: NetVLAD (Arandjelovic et al., 2016), SFRS (Ge et al., 2020), CosPlace (Berton et al., 2022a), MixVPR (Ali-Bey et al., 2023b), EigenPlaces (Berton et al., 2023), CricaVPR (Lu et al., 2024b), SALAD (Izquierdo and Civera, 2024), BOQ (Ali-bey et al., 2024), SALAD-CM (Izquierdo and Civera, 2025), SuperVLAD (Lu et al., 2024d), EMVP (Qiu et al., 2024), and the two re-ranking (two-stage) pipelines SelaVPR (Lu et al., 2024c) and FoL (Wang et al., 2025). Re-ranking pipelines add a computationally intensive local feature matching stage. EMVP and FoL are the current SOTA for single-stage global retrieval and two-stage VPR pipelines, respectively; implementation details for compared methods are given in App. A.4.
@@ 4.4 Ablation Study
Adding GWS to a SoftP configuration produces modest and unstable gains, suggesting sampling alone cannot exploit graph dynamics without Online Graph Creation. Enabling OGC and GWS while retaining CFP yields a notable improvement, for example Pitts30k-test R@1 of 94.9%, but still falls short of the configuration that includes SoftP. The full SAGE configuration achieves the best results, with R@1 of 95.4% on Pitts30k-test and R@1 of 98.0% on SPED, surpassing the baseline and intermediate variants. The greedy weighted clique expansion complements SoftP and OGC by focusing training on the most informative clusters and enhancing descriptor discriminability.
To evaluate the computational cost of our dynamic sampling strategy, we conducted an ablation study comparing the runtime and performance of online versus offline graph creation. As shown in Table 6, our online approach incurs a modest 17.7% increase in per-epoch training time.
@@
However, this modest overhead is a worthwhile investment, as it translates directly to superior accuracy, with the online SAGE model achieving higher recall on both SPED and MSLS-val.
@@
($d_{\text{model}},d_{\text{ff}}$) SPED Pitts30k
R@1 R@5 R@10 R@1 R@5 R@10
(512, 1024) 98.8 99.5 99.7 95.5 97.6 98.2
(768, 1536) 98.6 99.3 99.7 96.0 97.9 98.4
(768, 1024) 98.9 99.7 100 95.8 97.8 98.4
To demonstrate SAGE’s learning efficiency, we tracked its early-stage training performance on MSLS-val against a baseline using the Cliquemining (CM) strategy. As shown in Table 7, our model establishes a clear advantage by the fourth epoch (93.4% vs. 92.7% at R@1). The widening performance gap confirms that our dynamic sampling strategy fosters more efficient learning, leading to superior performance within the same number of epochs. As detailed in Tab. 8, we ablate the InteractHead’s internal dimensions. Setting $d_{model}=768$ and $d_{ff}=1024$ yields the best performance, notably achieving 98.9% R@1 on SPED. Expanding $d_{ff}$ to 1536 marginally improves Pitts30k but degrades SPED. Thus, we adopt the (768, 1024) configuration, balancing capacity and generalization.
@@ 5 Conclusion
In this paper, we presented SAGE, a unified framework that redefines Visual Place Recognition training by shifting from static sampling strategies to a dynamic, slow thinking paradigm. By synergizing the lightweight Soft Probing module with InteractHead, SAGE effectively amplifies fine-grained discriminative cues and models cross-image correlations, ensuring robust feature representation even under drastic appearance variations. Crucially, our novel Online Graph Creation and Greedy Weighted Sampling mechanisms ensure that the mining process continuously synchronizes with the evolving embedding space, allowing the model to relentlessly focus on the most informative geo-visual neighborhoods. Extensive evaluations across eight diverse benchmarks demonstrate that SAGE establishes a new SOTA, delivering exceptional retrieval accuracy while maintaining remarkable parameter efficiency through a frozen DINOv2 backbone. This provides a scalable and efficient foundation for future large-scale visual geo-localization systems.
6 Acknowledgments
This work was supported by the Beijing Natural Science Foundation (No.JQ23014), National Natural Science Foundation of China (No.62271074), Taishan Scholars Program (No.TSQN202507241), Key R&D Program of Shandong Province, China (No.2025KJHZ013), Shandong Provincial University Youth Innovation and Technology Support Program (No.2022KJ291), Shandong Provincial Natural Science Foundation for Young Scholars Program (No.ZR2025QC1627), and Qilu University of Technology (Shandong Academy of Sciences) Youth Outstanding Talent Program (No.2024QZJH02).
@@ A.1 Visualization of results
Figure 8: Visualization of SoftP’s learned feature importance. The visualizations demonstrate that SoftP automatically learns to focus on stable, fine-grained landmarks (e.g., building facades, structural details) while effectively ignoring non-discriminative regions (sky, road) and transient objects (vehicles, pedestrians).
@@ A.2 Additional Results
Building on this, Tab. 10 extends the evaluation to more demanding datasets characterized by severe domain shifts, including AmsterTime (historical vs. modern), Tokyo24/7 (extreme viewpoint and time-of-day changes), Pitts250k-test (large-scale urban scenes), and Eynsham (rural route). The performance trend remains robust, with accuracy scaling gracefully with descriptor dimensionality.
The improvement is particularly pronounced on AmsterTime, where the R@1 score surges from 23.4% at 128-D to 83.5% at 8448-D. This demonstrates SAGE’s exceptional capability to handle extreme appearance variations, validating the effectiveness of our proposed dynamic geo-visual graph exploration and feature enhancement strategies. To further assess the generalization capabilities of our method, we conducted experiments on the SF-small benchmark.
@@
Derived from SF-XL (Berton et al., 2022a), this dataset is particularly well-suited for evaluating generalization as its queries are drawn from distinct, non-adjacent locations.
@@
To further validate the effectiveness of SAGE, we present an extended comparison against several recently published SOTA methods from top-tier venues. These include VLAD-BuFF (Khaliq et al., 2024), EDTFormer (Jin et al., 2025b), and EffoVPR (Tzachor et al., 2025). The results, shown in Table 13, demonstrate that SAGE consistently outperforms these strong baselines. SAGE achieves superior performance across the majority of benchmarks. Although EffoVPR obtains a slightly higher R@1 on Tokyo24/7, it is a two-stage reranking method, which introduces additional computational overhead.
@@ A.3 Dataset Details
Pitts30k-test (Torii et al., 2013). The Pitts30k-test dataset is a subset of the Pittsburgh 250k dataset collected from Google Street View panoramas with GPS labels. It contains urban street-view images from Pittsburgh, Pennsylvania, USA, covering diverse city environments such as roads, bridges, and buildings, with large variations in viewpoint, season, and illumination. Each location provides multiple viewpoint images, and the dataset is primarily used for evaluating VPR models in urban scenarios.
@@
SPED (Chen et al., 2017). The SPED dataset, also known as the Specific PlacEs Dataset, consists of images captured by fixed surveillance cameras over extended time periods, covering significant changes in illumination, weather, and seasons. It contains unique locations, each with hundreds of images taken at different times of day and year, providing a challenging benchmark for long-term VPR under extreme appearance variations.
@@
MSLS-val (Warburg et al., 2020). The MSLS-val dataset is the validation split of the Mapillary Street-Level Sequences dataset, which contains street-level imagery sequences from cities worldwide captured with various devices such as smartphones, dashcams, and professional mapping cameras. The validation set covers multiple cities, seasons, and weather conditions, and is widely used for tuning and evaluating VPR models under diverse geographic and environmental variations.
@@
Nordland (Sünderhauf et al., 2013). The Nordland dataset consists of front-facing video recordings from a train journey along the same railway route in Norway, captured in all four seasons: spring, summer, autumn, and winter. The route and camera viewpoints are fixed, ensuring identical spatial structure while presenting extreme appearance changes due solely to seasonal and weather differences, making it a valuable benchmark for cross-season VPR research.
@@
Tokyo247 (Torii et al., 2015). The Tokyo247 dataset contains urban street-view images from Tokyo, Japan, primarily sourced from Google Street View panoramas with GPS labels, and is named for including both day and night imagery to represent long-term scene changes. It features dense metropolitan areas with tall buildings as well as some open spaces, offering large variations in viewpoint, illumination, and dynamic objects, and is used to evaluate VPR models in complex urban environments.
@@
AmsterTime (Yildiz et al., 2022). The AmsterTime dataset is a time-lapse street-view dataset of Amsterdam, containing images of the same locations captured in different years, such as 2008 and 2014, using Google Street View panoramas. It reflects long-term changes in building facades and urban infrastructure, providing a benchmark for evaluating VPR models under temporal urban transformations.
@@
Pitts250k-test (Torii et al., 2013). The Pitts250k-test dataset is the test split of the Pittsburgh 250k dataset, consisting of street-view images from Pittsburgh and surrounding areas, captured from Google Street View panoramas with GPS labels. It offers diverse urban scenes with dense road networks, buildings, and bridges, and serves as a large-scale benchmark for evaluating the scalability and robustness of VPR systems.
@@
Eynsham (Cummins and Newman, 2010; Berton et al., 2022b). The Eynsham dataset is a GPS-synchronized street-view image sequence collected along approximately 35 km of driving routes in Eynsham, Oxfordshire, UK, using a vehicle-mounted camera. It provides continuous video frames with precise ground-truth positions, making it suitable for sequence-based place recognition and loop closure detection experiments.
@@
SF-XL (Berton et al., 2022a). The San Francisco eXtra Large (SF-XL) dataset is a massive and dense new benchmark designed to push visual geo-localization research towards realistic, city-wide applications. Comprising over 41 million images captured across a decade, it presents significant real-world challenges such as long-term temporal variations and a domain shift between its Street View database and crowd-sourced queries.
@@ A.4 Compared Methods Details
NetVLAD (Arandjelovic et al., 2016) [^1] https://github.com/Nanne/pytorch-NetVlad. A classic VPR method with a learnable VLAD layer pluggable into any CNN. Uses VGG-16 backbone, trained on Pitts30k, optimized via weakly supervised ranking loss, outperforming traditional methods.
@@
SFRS (Ge et al., 2020) [^2] https://github.com/yxgeee/OpenIBL. Addresses GPS noise by mining hard positives via self-supervised fine-grained region similarities, multi-generation training. Based on NetVLAD, VGG-16 backbone, trained on Pitts30k, outperforming state-of-the-art then.
@@
CosPlace (Berton et al., 2022a) [^3] https://github.com/gmberton/CosPlace. Solves scalability in large-scale localization by framing training as classification. Constructs SF-XL dataset, uses CosPlace Groups. VGG-16/ResNet backbone, outputs 512D descriptors, low memory usage, suitable for city-scale applications.
@@
MixVPR (Ali-Bey et al., 2023b) [^4] https://github.com/amaralibey/MixVPR. This is a novel holistic feature aggregation method for VPR. It takes feature maps from pre-trained backbones as global features and iteratively incorporates global relationships into each feature map through stacked Feature-Mixer blocks (composed solely of multi-layer perceptrons), without the need for local or pyramidal aggregation. Using backbones like ResNet and trained on datasets such as GSV-Cities, it outperforms existing methods by a large margin with less than half the number of parameters compared to CosPlace and NetVLAD.
@@
R2Former (Zhu et al., 2023) [^5] https://github.com/bytedance/R2Former. This method introduces R2Former, a unified framework that employs a pure Transformer architecture to handle both global retrieval and local reranking in a single, end-to-end model. Its novel reranking module replaces slow geometric verification with a learnable Transformer that analyzes richer cues like feature correlation and attention, achieving state-of-the-art accuracy while dramatically reducing inference time and memory usage.
@@
EigenPlaces (Berton et al., 2023) [^6] https: //github.com/gmberton/EigenPlaces. This method embeds viewpoint robustness into learned global descriptors by clustering training data to explicitly present the model with different views of the same points of interest, without extra supervision. Using backbones like VGG-16 or ResNet with GeM pooling and trained on the SF-XL dataset, it outperforms state-of-the-art methods on most datasets, requiring $60\%$ less GPU memory for training and using $50\%$ smaller descriptors.
@@
SelaVPR (Lu et al., 2024c) [^7] https://github.com/Lu-Feng/SelaVPR. It proposes a hybrid global-local adaptation method that adapts pre-trained foundation models (e.g., DINOv2) via lightweight adapters without modifying the pre-trained model parameters, efficiently generating global features for candidate retrieval and local features for re-ranking. It also introduces a mutual nearest neighbor local feature loss to avoid time-consuming spatial verification. Outperforming state-of-the-art methods on benchmarks like MSLS, it consumes only about $3\%$ of the retrieval time of RANSAC-based two-stage methods.
@@
CricaVPR (Lu et al., 2024b) [^8] https://github.com/Lu-Feng/CricaVPR. It proposes cross-image correlation-aware representation learning, using attention to correlate features of multiple images in a batch, enabling each image feature to gain useful information from others for enhanced robustness. A multi-scale convolution-enhanced adaptation method is designed to insert lightweight adapters into frozen pre-trained foundation models (e.g., DINOv2) to introduce multi-scale local information. It outperforms state-of-the-art methods by a large margin on multiple benchmarks with shorter training time.
@@
SALAD (Izquierdo and Civera, 2024) [^9] https://github.com/serizba/salad. It reformulates NetVLAD’s soft assignment of local features to clusters as an optimal transport problem, considering both feature-to-cluster and cluster-to-feature relations, and introduces a ’dustbin’ cluster to discard non-informative features (Xu et al., 2024). Using DINOv2 as the backbone with fine-tuning, it trains in only 4 epochs. This single-stage method outperforms both single-stage and two-stage methods, with fast inference speed.
@@
BoQ (Ali-bey et al., 2024) [^10] https://github.com/amaralibey/ Bag-of-Queries. This method learns a set of global queries and uses cross-attention to probe input features for consistent information aggregation. It supports both CNN and Vision Transformer backbones, trained on the GSV-Cities dataset. As a one-stage global retrieval method without re-ranking, surpasses two-stage methods, and is fast and efficient.
@@
SALAD-CM (Izquierdo and Civera, 2025) [^11] https://github.com/serizba/cliquemining.This work addresses the insufficient Geographic Distance Sensitivity (GDS) of existing VPR models by proposing a novel sample mining strategy. It constructs a graph of visually similar images and samples cliques (sets of geographically close images) from the graph as training batches to enhance the model’s ability to distinguish small-range geographic distances. Based on models like DINOv2 SALAD and MixVPR, trained on densely sampled datasets such as MSLS and Nordland, it significantly improves recall without increasing inference computational overhead.
@@
SuperVLAD (Lu et al., 2024d) [^12] https://github.com/lu-feng/SuperVLAD. This method improves NetVLAD by removing cluster centers and using a small number of clusters , enhancing cross-domain generalization and simplifying the model. It also proposes 1-Cluster VLAD, which generates extremely low-dimensional descriptors by introducing “ghost clusters” and outperforms methods like GeM pooling with the same dimension. Using Transformer backbones (e.g., DINOv2) and trained on datasets like Pitts30k, it outperforms existing methods with lower feature dimensions.
@@
EMVP (Qiu et al., 2024) [^13] https://github.com/vincentqqb/EMVP. This method leverages Visual Foundation Models(e.g., DINOv2) and proposes a Centroid-Free Probing (CFP) stage that uses second-order features to better adapt VFM descriptors. It introduces a Dynamic Power Normalization (DPN) module to adaptively preserve task-specific information in both recalibration and CFP stages, forming a Parameter Efficiency Fine-Tuning (PEFT) pipeline. It achieves excellent performance on datasets like MSLS and Pitts250k, saving $64.3\%$ trainable parameters compared to existing state-of-the-art PEFT methods.
@@
FoL (Wang et al., 2025) [^14] https://github.com/chenshunpeng/FoL. This two-stage VPR method models reliable discriminative regions via Extraction-Aggregation Spatial Alignment Loss (SAL) and Foreground-Background Contrast Enhancement Loss (CEL), guiding global feature generation and efficient re-ranking. It introduces a weakly supervised local feature training strategy based on pseudo-correspondences and a discriminative region-guided efficient re-ranking pipeline. Using DINOv2 as the backbone and trained on GSV-Cities, it outperforms existing two-stage methods on multiple benchmarks with higher computational efficiency.
@@
VLAD-BuFF (Khaliq et al., 2024) [^15] https://github.com/Ahmedest61/VLAD-BuFF. This method improves visual place recognition accuracy by implementing a burst-aware weighting mechanism that discounts repetitive features to emphasize more distinctive visual cues. Simultaneously, it achieves high computational efficiency by using a pre-projection layer for local features, initialized with PCA, which enables rapid aggregation in a lower-dimensional space while maintaining high recall.